# Kramer’s Escape Rate and Phase Transition Dynamics in AdS Black Holes with Dark Structures

### v2

@@ 6 Conclusion
The study of black hole phase transitions based on the Gibbs free energy function has played a significant and influential role in advancing our understanding of black hole thermodynamics. However, a common issue with this conventional approach is that it seems to consider phase transitions only in a static form, leaving questions about the dynamics of phase transitions largely unanswered. The proposition that the dynamics of phase transitions in black holes are governed by thermodynamic and statistical forces has led to the possibility of examining phase transitions based on the stochastic Langevin equation or the Fokker-Planck equation [33], considering the ’mean first passage time’[34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. In this context, the introduction of the Kramers escape rate in the study of first-order phase transitions[60] prompted us to examine various free energy landscape forms and subsequently investigate the Kramers escape rate for the phase transitions of more complex black holes . These black holes are specifically formed based on cosmic and stringy features, meaning they interact more with the characteristics of dark structures. Our goal was to see what effect the addition of dark structure and stringy parameters would have on the dynamic behavior of phase transitions in this category of black holes.
Accordingly, in this paper, we have dynamically investigated the transition from a small black hole to a large black hole in first-order phase transitions from the perspective of the Kramers escape rate. Considering the escape rate as a function of distance or black hole radius, we examined whether the dynamic behavior of this phase transition for the black hole under study is exactly as expected at all moments of the transition. That is, firstly, the escape rate initially increases with the radius from a local minimum towards a global minimum and, after reaching a certain maximum, gradually decreases due to the activation of a reaction mechanism until it tends to zero. Secondly, the rate of escape and leakage from a small black hole to a large black hole always follows a logical and direct trend.
With this background, we studied the models (NLM-C-Q-PFD) and (EGB-YM-CS) and found that the escape rate diagram as a function of radius shows something different in specific regions. This means that the diagrams at the beginning, middle, and near the end of phase transition, at a specific radius, show equal probabilities of transition from a small to a large black hole and vice versa, which we have called contact coordinates in this paper. Beyond this radius, in a specific range up to the final point of phase transition, the reverse process becomes completely probable, meaning the probability of transitioning from a large black hole to a small black hole becomes dominant, and it seems that black holes prefer to travel from a large black hole to a small one. An interesting point about this region is that at the onset of phase transition, this area is very small, and as we approach the end in the sequence of frames, this area grows significantly.
Since, on the one hand, the growth of the probability of the reverse process corresponds to the increase in radius and, on the other hand, it is in the direction of the end of the phase transition process, it seems that in fact, this reverse trend is a reactive mechanism by the black hole. It is actually an effort to limit and control the phase transition to prevent excessive radial growth and prevent any uncontrolled instability so that the black hole can conclude the process towards a larger black hole under conditions that preserve the stability of the final structure and do not disrupt the survival of the black hole.
@@
7 Appendix A: Behavioral sequence of the free energy landscape for (EGB-YM-CS) black hole
@@
The swallow tail representation of the Gibbs free energy landscape in terms of t and the behavioral sequence of the potential function in terms of distance and temperature for the (EGB-YM-CS) black hole is shown in Fig (10), (11).

### v3

@@ VI Conclusion
The study of black hole phase transitions based on the Gibbs free energy function has played a significant and influential role in advancing our understanding of black hole thermodynamics. However, a common issue with this conventional approach is that it seems to consider phase transitions only in a static form, leaving questions about the dynamics of phase transitions largely unanswered. The proposition that the dynamics of phase transitions in black holes are governed by thermodynamic and statistical forces has led to the possibility of examining phase transitions based on the stochastic Langevin equation or the Fokker-Planck equation 21 , considering the ’mean first passage time’22 ; 23 ; 24 ; 25 ; 26 ; 27 ; 28 ; 29 ; 30 ; 31 ; 32 ; 33 ; 34 ; 35 ; 36 ; 37 ; 38 ; 39 ; 40 ; 41 ; 42 ; 43 ; 44 . In this context, the introduction of the Kramers escape rate in the study of first-order phase transitions48 prompted us to examine various free energy landscape forms and subsequently investigate the Kramers escape rate for the phase transitions of more complex black holes . These black holes are specifically formed based on cosmic and stringy features, meaning they interact more with the characteristics of dark structures. Our goal was to see what effect the addition of dark structure and stringy parameters would have on the dynamic behavior of phase transitions in this category of black holes. Accordingly, in this paper, we have dynamically investigated the transition from a small black hole to a large black hole in first-order phase transitions from the perspective of the Kramers escape rate. Considering the escape rate as a function of distance or black hole radius, we examined whether the dynamic behavior of this phase transition for the black hole under study is exactly as expected at all moments of the transition. That is, firstly, the escape rate initially increases with the radius from a local minimum towards a global minimum and, after reaching a certain maximum, gradually decreases due to the activation of a reaction mechanism until it tends to zero. Secondly, the rate of escape and leakage from a small black hole to a large black hole always follows a logical and direct trend. With this background, we studied the models (NLM-C-Q-PFD) and (EGB-YM-CS) and found that the escape rate diagram as a function of radius shows something different in specific regions. This means that the diagrams at the beginning, middle, and near the end of phase transition, at a specific radius, show equal probabilities of transition from a small to a large black hole and vice versa, which we have called contact coordinates in this paper. Beyond this radius, in a specific range up to the final point of phase transition, the reverse process becomes completely probable, meaning the probability of transitioning from a large black hole to a small black hole becomes dominant, and it seems that black holes prefer to travel from a large black hole to a small one. An interesting point about this region is that at the onset of phase transition, this area is very small, and as we approach the end in the sequence of frames, this area grows significantly. Since, on the one hand, the growth of the probability of the reverse process corresponds to the increase in radius and, on the other hand, it is in the direction of the end of the phase transition process, it seems that in fact, this reverse trend is a reactive mechanism by the black hole. It is actually an effort to limit and control the phase transition to prevent excessive radial growth and prevent any uncontrolled instability so that the black hole can conclude the process towards a larger black hole under conditions that preserve the stability of the final structure and do not disrupt the survival of the black hole.
@@
VII Appendix A: Behavioral sequence of the free energy landscape for (EGB-YM-CS) black hole
@@
The swallow tail representation of the Gibbs free energy landscape in terms of t and the behavioral sequence of the potential function in terms of distance and temperature for the (EGB-YM-CS) black hole is shown in Figs. (10), (11).